# Convergence Analysis and Acceleration of Fictitious Play for General Mean-Field Games via the Best Response

### v1

@@ 2.2.5 Potential MFGs
Remark 2.28 (Assumptions on $F$).
The potential $F$ being convex and $L_{f}$-Lipschitz smooth is equivalent to $f$ being monotone and $L_{f}$-Lipschitz continuous.
By coercivity of the objective function $j$, we can restrict our focus to a subset of $C$ with diameter $D<\infty$. With Lemma 2.27, we have the convergence for potential games.
@@
Under assumptions 1-2, if the sequence generated by (72)-(73) with $\delta_{k}=\frac{2}{k+2}$ is in a subset of $C$ with diameter $D$, then the sequence satisfies
@@
Remark 2.30 (Comparison with existing optimization methods).
When the Mean Field Game (MFG) is potential, other optimization algorithms have been employed to solve the variational formulation. In comparison to the proximal gradient approach [51], the stepsize and convergence properties of fictitious play are independent of the regularity of the non-smooth component, i.e. dynamic cost involving $\frac{m}{\rho}$. Consequently, this method is capable of effectively handling cases where $\rho$ is close to or equal to zero. The convergence of primal-dual [44] depends on the regularity of the constraint. For MFG, the constraint involves differential operators which are unbounded. When solving the discretized system, the condition number of the derivative operator grows to infinity as the mesh step decreases to 0 and the primal-dual will take more iterations to converge. Compared to primal-dual [44], our convergence is independent of the regularity of the constraint and therefore the mesh step. More details are in Remark 3.5. Compared to ADMM and augmented Lagrangian [9, 10], the subproblem in fictitious play can be solved with better accuracy in less time.
We emphasize that the sublinear convergence of the primal gap does not require the directional strong convexity of $r(\rho,m)$ in $\rho$. To close this section, we show that if $r$ is directional strongly convex between $(\rho,m)$ and its best response $(\hat{\rho},\hat{m})$, then the primal gap and gain of the best response (dual gap) are equivalent.
Lemma 2.31.
Under Assumption 2, if $r$ is $\lambda_{r}$-strongly convex between $(\rho,m)$ and its best response, $s$ is convex and $L_{f}$-Lipschitz smooth, then
Proof.
We first show that $h(\rho,m)\leq g(\rho,m)$.
$h(\rho,m)$ $=r(\rho,m)-r(\rho^{*},m^{*})+s(\rho)-s(\rho^{*})+\langle f_{T},\rho_{T}-\rho^{*}_{T}\rangle_{\Omega}$ (85)
$\leq r(\rho,m)-r(\rho^{*},m^{*})+\langle\nabla s(\rho),\rho-\rho^{*}\rangle_{Q_{T}}+\langle f_{T},\rho_{T}-\rho^{*}_{T}\rangle_{\Omega}$
$\leq g(\rho,m).$
Here the first and third lines are by definition and the second line is by convexity of $s$.
Next, we define
$(\hat{\rho},\hat{m})=\operatorname*{argmin}_{\tilde{\rho},\tilde{m}\in C}r(\tilde{\rho},\tilde{m})+\langle\nabla s(\rho),\tilde{\rho}\rangle_{Q_{T}}+\langle f_{T},\tilde{\rho}_{T}\rangle_{\Omega},$
and
$(\rho^{+},m^{+})=(1-\delta)({\rho},{m})+\delta(\hat{\rho},\hat{m}),$
where $\delta\in(0,1]$. Then repeating the analysis in (76), (77) and (78) gives us
$h(\rho^{+},m^{+})\leq h(\rho,m)-\delta g(\rho,m)+\delta^{2}\frac{L_{f}}{2}\|\hat{\rho}-\rho\|^{2}.$ (86)
Since $h\geq 0$ and by Lemma 2.11, we have
@@ 3 Discretization and Implementation
This section first discusses the finite difference approximation of (2) on $\Omega=[0,1]$. This discretization can be easily generalized to cuboid in high-dimension. Based on the discretization, we propose a fictitious play accelerated by employing back-tracking line search and hierarchical of grids for solving the discrete system.
Let $n_{t}$ be a positive integer and $\Delta t=\frac{T}{n_{t}},t_{n}=n\Delta t,n=0,\cdots,n_{t}$. Similarly, let $n_{x}$ be a positive integer and $\Delta x=\frac{1}{n_{x}},x_{i}=i\Delta x,i=0,\cdots,n_{x}$. Denote $\mathcal{G}$ as the collection of all grid points $(x_{i},t_{n}),i=0,\cdots,n_{x},n=0,\cdots,n_{t}$. For any function $u$ defined on $\Omega\times[0,T]$, we denote the approximation of $u$ on $\mathcal{G}$ as $u_{\mathcal{G}}$ where $(u_{\mathcal{G}})_{i,n}$ approximates $u(x_{i},t_{n})$. When the context provides no ambiguity, we omit $\mathcal{G}$ in the subscript. In this paper, we consider the homogeneous Neumann boundary condition. To tackle this in the discretization, we introduce ghost points $i=-1$ and $i=n_{x}+1$ and we set $u_{-1}=u_{0}$ and $u_{n_{x}+1}=u_{n_{x}}$. This results in the following finite difference operators on the grid $\mathcal{G}$. Firstly, the elementary one-side finite difference operators are defined as
@@
Then we define the central difference operator $(D_{x}^{c}u)_{i,n}$ as the average of the one-sided finite difference, and $[D_{x}u]_{i,n}$ as the collection of the one-sided finite differences at $(x_{i},t_{n})$
@@
The way we introduce the ghost point values makes the discrete Laplacian operator self-adjoint under the above inner product. To preserve the adjoint relation between gradient and negative divergence, for a given velocity field $v$ on $\mathcal{G}$, we define its divergence to be $-D_{x}^{*}[v]$ where $D_{x}^{*}$ is the adjoint operator of $D_{x}$,
@@
We consider Lax-Friedrichs Hamiltonian
@@
We approximate the interaction cost $f(x,\rho)$ and $f_{T}(x,\rho)$ at $x_{i}$ by $f_{\mathcal{G}}(x_{i},\rho_{\mathcal{G}})$ and $f_{T,\mathcal{G}}(x_{i},\rho_{\mathcal{G}})$, respectively. For any function $\rho_{\mathcal{G}}$ on grid $\mathcal{G}$, we define its corresponding piece-wise constant function on $\Omega$ as $\bar{\rho}_{\mathcal{G}}$ with $\bar{\rho}_{\mathcal{G}}(x)=\rho_{i},x_{i}-\frac{\Delta x}{2}\leq x<x_{i}+\frac{\Delta x}{2}$. For consistency, we assume that there exists a constant $C$ independent of $\Delta x$ and $\rho$ such that for any $x_{i}$ and $\rho_{\mathcal{G}}$,
@@
    If $f(x,\rho)=\int_{\Omega}K(x,y)\rho(y)\mathrm{d}y$ is a nonlocal cost defined by some kernel function $K$, it is usually approximated by $f_{\mathcal{G}}(x_{i},\rho_{\mathcal{G}})=\sum_{j=0}^{n_{x}}K(x_{i},x_{j})\rho_{j}\Delta x$
@@
    In section 4, we consider a cost of form $f(x,\rho)=f_{0}(x,\mu_{1}(\rho),\mu_{2}(\rho))$ where $\mu_{1}(\rho)=\int_{\Omega}x\rho(x)\mathrm{d}x$ is the first moment and $\mu_{2}(\rho)=\int_{\Omega}x^{2}\rho(x)\mathrm{d}x$ is the second moment of $\rho$. We first approximate the first and second moments with $\mu_{1,\mathcal{G}}(\rho_{\mathcal{G}})=\sum_{i=0}^{n_{x}}x_{i}\rho_{i}\Delta x$ and $\mu_{2,\mathcal{G}}(\rho_{\mathcal{G}})=\sum_{i=0}^{n_{x}}x_{i}^{2}\rho_{i}\Delta x$ and then define $f_{\mathcal{G}}(x_{i},\rho_{\mathcal{G}})=f_{0}(x_{i},\mu_{1,\mathcal{G}}(\rho_{\mathcal{G}}),\mu_{2,\mathcal{G}}(\rho_{\mathcal{G}}))$.
@@
According to [3, 2], 1-3 are needed to establish the existence of the solution to the discrete MFG system (90) and 1-4 are needed to establish the uniqueness of the solution. For the detail, we refer readers to [3, 2]. It is also established therein that the solution to the discrete system (90) converges to the solution of the continuous system (2) as the mesh step approaches zero.
@@
Requiring $R_{\text{FP}}(\rho,\phi)=\mathbf{0}$ is equivalent to applying a semi-implicit Lax-Friedriches scheme to the Fokker-Planck equation with zero-flux boundary condition, i.e.
@@
The discrete MFG system (90) is semi-implicit. To be precise, the discrete HJB (88) is implicit in $\phi$ and explicit in $\rho$ while the discrete FP (89) is implicit in $\rho$ and explicit in $\phi$. While one can also choose the implicit scheme, the semi-implicit scheme preserves the time-adjoint relation between $\rho$ and $\phi$, which is essential for preserving the variational formulation and equilibrium formulation in discretization. Remark 3.4 and 3.5 give more concrete explanations. In addition to the above advantage, according to [3], it is easier to show the uniqueness of the solution to the discrete MFG system with the semi-implicit scheme.
@@
being the discrete constraint set and $\phi_{i,n}$ being the dual variable to the $(i,n)$-th constraint of $\rho,m$. Such variational formulation does not exist if we discretize HJB and FP with implicit schemes.
@@ 3.1 Fictitious play for the discrete system
While requiring $R_{\text{HJB}}(\rho,\phi)=\mathbf{0}$ or $R_{\text{FP}}(\rho,\phi)=\mathbf{0}$ can be addressed backward or forward with relative ease, satisfying both simultaneously is more complex. There is no straightforward time-marching method to achieve this; consequently, the entire system must be solved concurrently. This is challenging due to the presence of nonlinear Hamiltonian and cost terms. To numerically solve the discrete system (90) we alternatively solve the HJB (88) and the FP (89) equations. Algorithm 3 summarizes the procedure. In practice, we solve the HJB backward with Newton’s method.
@@
Remark 3.5 (Equilibrium form for the discrete system 90 and reformulation of Algorithm 3).
@@
$(\hat{\rho}^{(k)},\hat{m}^{(k)})\in\operatorname*{argmin}_{(\hat{\rho},\hat{m})\in C_{\mathcal{G}}}J_{\mathcal{G}}(\hat{\rho},\hat{m};\rho^{(k-1)}),$ (94)
@@
This is the discrete counterpart of (16)-(17) for the continuous fictitious play (Algorithm 1). Therefore, the convergence analysis for Algorithm 1 also holds for Algorithm 3.
It is important to notice that the convergence rate depends only on the Lipschitz constant $L$ and convexity $\mu$ which are $\mathcal{O}(1)$ with respect to $\Delta x,\Delta t$. Therefore, the convergence rate of the discrete fictitious play (Algorithm 3) does not depend on the mesh step. This is a big advantage compared to primal-dual since the convergence rate of primal-dual depends on the operator norm of constraint and in our case, the differential operator norm is $\mathcal{O}(\frac{1}{\min\{\Delta t,\Delta x\}})$.
@@ 3.2 Acceleration: incorporating hierarchical grids
Our method uses a hierarchy of grids but is different from multigrid methods [48] used in [7, 8, 13], where they have both prolongation and restriction. Essentially, our hierarchical grid approach captures low-frequency modes on coarse grids first and gradually adds the high-frequency modes on fine grids. On the contrary, multigrid methods first capture the high-frequency modes on fine grids and reduce the low-frequency error on coarse grids. The idea of our hierarchy grid acceleration is similar to the multilevel approach in [51]. But since their algorithm couples $\rho$ and $m$, their interpolation is acted on both variables, while we only interpolate $\rho$ or $\phi$. In addition, our scheme is more stable than [51] when $\rho$ is compactly supported or is close to 0 in the domain.
@@ 4 Numerical Results
In this section, we conduct various experiments to show the effectiveness and efficiency of our hierarchical grid fictitious play algorithm. All of our numerical experiments are implemented in Matlab on a PC with an Apple M3 Pro chip and 18 GB memory.
@@
While theoretically $g^{k}\geq 0$, numerically it may become negative when $g^{(k)}$ is close to 0. Therefore we use $|g^{(k)}|\leq\epsilon$ as the stopping criteria and when presenting the semilogy plot of $g^{(k)}$, we use $|g^{(k)}|$. The second is the consecutive residue $\|\hat{\rho}^{(k)}-\rho^{(k-1)}\|_{\mathcal{G}}$. As discussed in Section 2.2, under certain conditions, the distance between $\rho^{(k-1)}$ and the Nash Equilibrium $\rho^{*}$ is bounded by the consecutive residue. The third measurement is FP residue $\left\|R_{\text{FP}}\left(\rho^{(k)},\hat{\phi}^{(k+1)}\right)\right\|_{\mathcal{G}}.$ Since $\hat{\phi}^{(k+1)}$ is obtained from solving HJB whose source term is given by $\rho^{(k)}$, the HJB residue $\left\|R_{\text{HJB}}\left(\rho^{(k)},\hat{\phi}^{(k+1)}\right)\right\|_{\mathcal{G}}$ is only due to the discretization and HJB solver. The FP residue reflects how well $\rho^{(k)}$ approaches the fixed point of the best response mapping and should be small when $\rho^{(k)}$ is close to a Nash Equilibrium.
While the theoretical convergence result is established only for monotone games and density-independent terminal cost, the convergence analysis suggests that provided that the terms due to non-monotone cost and dependent $f_{T}$ are adequately controlled, the fictitious play still produces a decreasing sequence of $g^{(k)}$ and thus converges. Our numerical experiments also demonstrate success in non-monotone MFGs and more general forms of $f_{T}$.

### v2

@@ 3.6 Potential MFGs
By coercivity of the objective function $J$, we can restrict our focus to a subset of $C$ with diameter $D<\infty$. According to braun2023gcg-cgbook , $\delta_{k}=\frac{2}{k+2}$ is known as the open-loop step-size and is a standard choice in Frank-Wolfe when the information of the function is not available. Taking $\delta_{k}=\frac{2}{k+2}$ in Lemma 3.26, we have the convergence for potential games. We emphasize that the sublinear convergence of the primal gap does not require the directional strong convexity of $R(\rho,m)$ in $\rho$.
@@
Under assumptions 3.5, 3.6 and 3.24, if $F$ is convex on $\mathcal{P}(\mathbb{T}^{d})$ under $L^{2}$ metric and suppose that the sequence generated by fictitious play (24)–(5) with step size $\delta_{k}=\frac{2}{k+2}$ remains within a subset of $C$ with diameter $D$. Then the sequence satisfies the bound
@@
When the Mean Field Game (MFG) is potential, other optimization algorithms have been employed to solve the variational formulation. The convergence of primal-dual liu2021nonlocal depends on the regularity of the constraint. For MFG, the constraint involves differential operators that are unbounded. When solving the discretized system, the condition number of the derivative operator grows to infinity as the mesh step decreases to 0 and the primal-dual will take more iterations to converge. Compared to primal-dual liu2021nonlocal , our convergence is independent of the regularity of the constraint and therefore the mesh step. More details are in Remark 4.5. Compared to ADMM and augmented Lagrangian benamou2015augmented ; benamou2017variational , the subproblem in fictitious play can be solved with better accuracy in less time.
@@ 4 Discretization and Implementation
This section first discusses the finite difference approximation of (2) on $[0,1]$. This discretization can be easily generalized to a cuboid in high dimensions. Based on the discretization, we propose a fictitious play accelerated by employing a hierarchical grid for solving the discrete system.
Let $n_{t}$ be a positive integer and $\Delta t=\frac{T}{n_{t}},t_{n}=n\Delta t,n=0,\cdots,n_{t}$. Similarly, let $n_{x}$ be a positive integer and $\Delta x=\frac{1}{n_{x}},x_{i}=i\Delta x,i=0,\cdots,n_{x}$. Denote $\mathcal{G}$ as the collection of all grid points $(x_{i},t_{n}),i=0,\cdots,n_{x},n=0,\cdots,n_{t}$. For any function $u$ defined on $\mathbb{T}^{d}\times[0,T]$, we denote the approximation of $u$ on $\mathcal{G}$ as $u_{\mathcal{G}}$ where $(u_{\mathcal{G}})_{i,n}$ approximates $u(x_{i},t_{n})$. When the context provides no ambiguity, we omit $\mathcal{G}$. In this section, we consider the homogeneous Neumann boundary condition for the HJB equation and the zero-flux boundary condition for the FP equation. To tackle this in the discretization, we introduce the following finite difference operators on the grid $\mathcal{G}$. Firstly, the elementary one-sided finite difference operators are defined as
@@
Then we define the central difference operator $(D_{x}^{c}u)_{i,n}$ as the average of the one-sided finite difference, $[D_{x}u]_{i,n}$ as the collection of the one-sided finite differences at $(x_{i},t_{n})$
@@
The discrete Laplacian operator is self-adjoint under the above inner product. To preserve the adjoint relation between gradient and negative divergence, for a given velocity field $v$ on $\mathcal{G}$, we define its divergence to be $-D_{x}^{*}[v]$ where $D_{x}^{*}$ is the adjoint operator of $D_{x}$,
@@
We consider the Lax-Friedrichs Hamiltonian
@@
We approximate the interaction cost $f(x,\rho)$ and $f_{T}(x,\rho)$ at $x_{i}$ by $f_{\mathcal{G}}(x_{i},\rho_{\mathcal{G}})$ and $f_{T,\mathcal{G}}(x_{i},\rho_{\mathcal{G}})$, respectively. For any function $\rho_{\mathcal{G}}$ on grid $\mathcal{G}$, we define its corresponding piece-wise constant function on $\mathbb{T}^{d}$ as $\bar{\rho}_{\mathcal{G}}$ with $\bar{\rho}_{\mathcal{G}}(x)=\rho_{i},x_{i}-\frac{\Delta x}{2}\leq x<x_{i}+\frac{\Delta x}{2}$. For consistency, we assume that there exists a constant $C$ independent of $\Delta x$ and $\rho$ such that for any $x_{i}$ and $\rho_{\mathcal{G}}$,
@@
    If $f(x,\rho)=\int_{0}^{1}K(x,y)\rho(y)\mathrm{d}y$ is a nonlocal cost defined by some kernel function $K$, it is usually approximated by $f_{\mathcal{G}}(x_{i},\rho_{\mathcal{G}})=\sum_{j=0}^{n_{x}}K(x_{i},x_{j})\rho_{j}\Delta x$
@@
    In section 5, we consider a cost of form $f(x,\rho)=f_{0}(x,\mu_{1}(\rho),\mu_{2}(\rho))$ where $\mu_{1}(\rho)=\int_{0}^{1}x\rho(x)\mathrm{d}x$ is the first moment and $\mu_{2}(\rho)=\int_{0}^{1}x^{2}\rho(x)\mathrm{d}x$ is the second moment of $\rho$. We first approximate the first and second moments with $\mu_{1,\mathcal{G}}(\rho_{\mathcal{G}})=\sum_{i=0}^{n_{x}}x_{i}\rho_{i}\Delta x$ and $\mu_{2,\mathcal{G}}(\rho_{\mathcal{G}})=\sum_{i=0}^{n_{x}}x_{i}^{2}\rho_{i}\Delta x$ and then define $f_{\mathcal{G}}(x_{i},\rho_{\mathcal{G}})=f_{0}(x_{i},\mu_{1,\mathcal{G}}(\rho_{\mathcal{G}}),\mu_{2,\mathcal{G}}(\rho_{\mathcal{G}}))$.
@@
According to achdou2010mean ; Achdou2013fdcvg , 1-3 are needed to establish the existence of the solution to the discrete MFG system (102) and 1-4 are needed to establish the uniqueness of the solution. For the details, we refer readers to achdou2010mean ; Achdou2013fdcvg . It is also established therein that the solution to the discrete system (102) converges to the solution of the continuous system (2) as the mesh step approaches zero.
@@
Requiring $R_{\text{FP}}(\rho,\phi)=\mathbf{0}$ is equivalent to applying a semi-implicit Lax-Friedrichs scheme to the Fokker-Planck equation with zero-flux boundary condition, i.e.
@@
The discrete MFG system (102) is semi-implicit. To be precise, the discrete HJB (100) is implicit in $\phi$ and explicit in $\rho$ while the discrete FP (101) is implicit in $\rho$ and explicit in $\phi$. While one can also choose the implicit scheme, the semi-implicit scheme preserves the time-adjoint relation between $\rho$ and $\phi$, which is essential for preserving the variational formulation and equilibrium formulation in discretization. Remark 4.4 and 4.5 give more concrete explanations. In addition to the above advantage, according to achdou2010mean , it is easier to show the uniqueness of the solution to the discrete MFG system with the semi-implicit scheme.
@@
being the discrete constraint set and $\phi_{i,n}$ being the dual variable to the $(i,n)$-th constraint of $\rho,m$. Such a variational formulation does not exist if we discretize HJB and FP with implicit schemes.
@@ 4.1 Fictitious play for the discrete system
While requiring $R_{\text{HJB}}(\rho,\phi)=\mathbf{0}$ or $R_{\text{FP}}(\rho,\phi)=\mathbf{0}$ can be addressed backward or forward with relative ease, satisfying both simultaneously is more complex. There is no straightforward time-marching method to achieve this; consequently, the entire system must be solved concurrently. This is challenging due to the presence of nonlinear Hamiltonian and cost terms. To numerically solve the discrete system (102), we alternatively solve the HJB (100) and the FP (101) equations. Algorithm 3 summarizes the procedure. In practice, we solve the HJB backward with Newton’s method.
@@
Remark 4.5 (Equilibrium form for the discrete system (102) and reformulation of Algorithm 3).
@@
$(\hat{\rho}^{(k)},\hat{m}^{(k)})\in\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C_{\mathcal{G}}}J_{\mathcal{G}}(\tilde{\rho},\tilde{m};\rho^{(k-1)}),$ (106)
@@
This is the discrete counterpart of (24)-(25) for the continuous fictitious play (Algorithm 1). Therefore, the convergence analysis for Algorithm 1 also holds for Algorithm 3 under sufficient regularity assumptions and fine meshes.
It is important to notice that the convergence rate depends only on the Lipschitz constant $L_{f}$ and convexity $\lambda_{R}$, which are $\mathcal{O}(1)$ with respect to $\Delta x,\Delta t$. Therefore, the convergence rate of the discrete fictitious play (Algorithm 3) does not depend on the mesh step. This is a big advantage compared to primal-dual since the convergence rate of primal-dual depends on the operator norm of constraint and in our case, the differential operator norm is $\mathcal{O}(\frac{1}{\min\{\Delta t,\Delta x\}})$.
@@ 4.2 Acceleration: incorporating hierarchical grids
Our method uses a hierarchy of grids but is different from multigrid methods Trottenberg2000multigrid used in achdou2012multigrid ; andreev2017multigridadmm ; briceno2018multigridadmm , where they have both prolongation and restriction. Essentially, our hierarchical grid approach captures low-frequency modes on coarse grids first and gradually adds the high-frequency modes on fine grids. On the contrary, multigrid methods first capture the high-frequency modes on fine grids and reduce the low-frequency error on coarse grids. The idea of our hierarchy grid acceleration is similar to the multilevel approach in yu2024pgdeucmfg . But since their algorithm couples $\rho$ and $m$, their interpolation acts on both variables, while we only interpolate $\rho$ or $\phi$.
@@ 5 Numerical Results
In this section, we conduct various experiments to show the effectiveness and efficiency of our hierarchical grid fictitious play algorithm. All of our numerical experiments are implemented in MATLAB on a PC with an Apple M3 Pro chip and 18 GB of memory.
@@
While theoretically $g^{k}\geq 0$, numerically it may become negative when $g^{(k)}$ is close to 0. Therefore, we use $|g^{(k)}|\leq\epsilon$ as the stopping criterion and when presenting the semilogy plot of $g^{(k)}$, we use $|g^{(k)}|$. The second is the consecutive residue $\|\hat{\rho}^{(k)}-\rho^{(k-1)}\|_{\mathcal{G}}$. As discussed in Section 3, under certain conditions, the distance between $\rho^{(k-1)}$ and the Nash Equilibrium $\rho^{*}$ is bounded by the consecutive residue. The third measurement is FP residue $\left\|R_{\text{FP}}\left(\rho^{(k)},\hat{\phi}^{(k+1)}\right)\right\|_{\mathcal{G}}.$ Since $\hat{\phi}^{(k+1)}$ is obtained from solving HJB whose source term is given by $\rho^{(k)}$, the HJB residue $\left\|R_{\text{HJB}}\left(\rho^{(k)},\hat{\phi}^{(k+1)}\right)\right\|_{\mathcal{G}}$ is only due to the discretization and HJB solver. The FP residue reflects how well $\rho^{(k)}$ approaches the fixed point of the best response mapping and should be small when $\rho^{(k)}$ is close to a Nash Equilibrium.
While the theoretical convergence result is established only for monotone games and density-independent terminal cost, the convergence analysis suggests that, provided that the terms due to non-monotone cost and dependent $f_{T}$ are adequately controlled, the fictitious play still produces a decreasing sequence of $g^{(k)}$ and thus converges. Our numerical experiments also demonstrate success in non-monotone MFGs and more general forms of $f_{T}$.
5.1 Divergence of fixed-point iteration
We first show that fixed-point iteration may diverge under a very simple and ideal setting. On $\mathbb{T}$, consider a MFG with $H(x,p)=\frac{1}{2}|p|^{2},f(x,\rho)=\rho(x)+\exp(\sin(2\pi x))$ and $f_{T}(x,\rho)=0$. This is a potential MFG with a monotone interaction cost and therefore admits at most one Nash Equilibrium. Let $\rho_{0}(x)=\rho_{G}(x;0,0.2)$. We first apply fictitious play to solve the problem, and the algorithm converges in 38 iterations, reaching a residue of $10^{-13}$. This gives us a numerical equilibrium $(\rho^{*},v^{*})$ up to machine accuracy. We then initialize the fixed-point iteration with the equilibrium and report the results in Figure 2. The plots show that the fixed-point iteration diverges, even starting from the Nash Equilibrium. In fact, it oscillates between two states in the end.
Figure 2: Numerical results of Example 5.1. The fixed-point algorithm diverges even starting from the Nash Equilibrium. Left: the semilog plot of $g^{(k)}$. Center: consecutive residue $\|\hat{\rho}^{(k)}-\rho^{(k-1)}\|_{\mathcal{G}}$. Right: the error $\|\rho^{(k)}-\rho^{*}\|_{\mathcal{G}}$.